# Normalize Filters! Classical Wisdom for Deep Vision

### v2

@@ 3.2 Filter normalization
If all filters in a layer are unnormalized, each response is scaled by the filter’s total weight sum, $\|w\|r(w)$, thereby altering the feature representation at that layer.
@@
$y$ $=\sum_{i=1}^{k}w_{i}x_{i}+b=\|w\|\sum_{i=1}^{k}\frac{w_{i}}{\|w\|}x_{i}+b={\color[rgb]{0,1,1}\|w\|}\sum_{i=1}^{k}\left({\color[rgb]{1,0,0}\frac{w^{+}_{i}}{\|w\|}-\frac{w^{-}_{i}}{\|w\|}}\right)x_{i}+{\color[rgb]{0,1,1}b}$ (12)
$\hskip 40.0pt\to{\color[rgb]{0,1,1}\underbrace{a}_{\text{scaling}}}\sum_{i=1}^{k}({\color[rgb]{1,0,0}\underbrace{\frac{w^{+}_{i}}{\|w^{+}\|+\varepsilon}-\frac{w^{-}_{i}}{\|w^{-}\|+\varepsilon}}_{\text{filter normalization}}})x_{i}+{\color[rgb]{0,1,1}\underbrace{b}_{\text{shifting}}},$ (13)
where $\varepsilon$ is a small constant (e.g., $10^{-6}$) to ensure numerical stability, and it is then followed by learnable scaling $a$ and shifting $b$, which model response weighting $\|w\|$ and offset $b$, respectively.
This normalization step enforces the filter to become either averaging if $w$ is all positive, or differencing if $w$ has both positive and negative weights, with predictable responses to intensity changes:
@@
    Averaging filters respond linearly to both gain and offset: $f(g\cdot x+o)=g\cdot f(x)+o$, where $\|w\|=1$ and $r(w)=1$.
@@ 4.3 Analysis and ablation studies
We identify unnormalized filters as a key deep learning limitation and propose filter normalization, ensuring atmospheric equivariance. We achieve consistent gains across CNNs and ViTs, outperforming larger models like CLIP, while enhancing robustness and generalization. Our code will be publicly available upon publication.

### v3

@@ 3.2 Filter normalization
If all filters in a layer are unnormalized, each response is scaled by the filter’s total weight sum, $\|w\|r(w)$, thereby altering the feature representation at that layer. However, if all filters are normalized, their responses are scaled uniformly by the same gain factor $g$, that is, they produce consistent and interpretable responses with respect to intensity changes.
@@
$y$ $=\sum_{i=1}^{k}w_{i}x_{i}+b=\|w\|\sum_{i=1}^{k}\frac{w_{i}}{\|w\|}x_{i}+b={\color[rgb]{1,0,1}\|w\|}\sum_{i=1}^{k}\left({\color[rgb]{0,0,1}\frac{w^{+}_{i}}{\|w\|}-\frac{w^{-}_{i}}{\|w\|}}\right)x_{i}+{\color[rgb]{1,0,1}b}$ (12)
$\hskip 40.0pt\to{\color[rgb]{1,0,1}\underbrace{a}_{\text{scaling}}}\sum_{i=1}^{k}{\color[rgb]{0,0,1}\underbrace{\left(\frac{w^{+}_{i}}{\|w^{+}\|+\varepsilon}-\frac{w^{-}_{i}}{\|w^{-}\|+\varepsilon}\right)}_{{\color[rgb]{0,0,1}\text{filter normalization}}}}x_{i}+{\color[rgb]{1,0,1}\underbrace{b}_{\text{shifting}}},$ (13)
where $\varepsilon$ is a small constant (e.g., $10^{-6}$) to ensure numerical stability. This is followed by learnable scaling $a$ and shifting $b$, which model response weighting $\|w\|$ and offset $b$, respectively.
This normalization step enforces the filter to become either averaging if $w$ is all positive, or differencing if $w$ has both positive and negative weights:
@@
    Averaging filters are equivariant to both gain and offset: $f(g\cdot x+o)=g\cdot f(x)+o$, where $\|w\|=1$ and $r(w)=1$.
@@ 4.3 Analysis and ablation studies
We identify unnormalized filters as a key deep learning limitation and propose filter normalization, ensuring atmospheric equivariance. We achieve consistent gains across CNNs and ViTs, outperforming larger models like CLIP, while enhancing robustness and generalization. Code available at: https://github.com/gperezs/normalized-convolutions.
@@ Acknowledgments
Normalize Filters! Classical Wisdom for Deep Vision